i. KYC company makes 27 orders of chairs in a year and orders 18,306 chairs each time. The fixed order costs are $424.89 per order and the carrying cost per unit is $59.18. The chairs are sold out before they are restocked. What are the shortage costs if the company orders the optimal quantity (Economic Order Quantity)?

Enter your answer rounded off to two decimal points. Do not enter comma or $ in the answer box.

ii. Deploy ABC Company sells 36,993 chairs a year at an average price per chair of $6.42. The carrying cost per unit is $16.13. The company orders 111 chairs at a time and has a fixed order cost of $182.79 per order. The chairs are sold out before they are restocked. What are the total costs if the company orders the optimal quantity (Economic Order Quantity)?

Enter your answer rounded off to two decimal points. Do not enter comma or $ in the answer box.
Answer #1

Solution :-

Total Shortage Costs = Fixed Costs * T/Q

Fixed Costs = 424.89

T =Total units sold= 27*18306 = 494262

Q =Total ordered = 27

Total Shortage Costs = 424.89 * 494262/18306 =

= $11472.03

As per HomeworkLib policy we need to answer only one question at once so please ask other as seperate one.